Given P(A) = 0.82, P(B) = 0.87, and the fact that events A and B are independent, What is P(A and B)?
Answer by ikleyn(53751) 
 .
Since events A and B are independent, it means, BY THE DEFINITION, that
the probability of the event (A and B) is the product P(A) and P(B)
P(A and B) = P(A)*P(B).
Hence, all you need to do is to multiply the two given numbers.
ANSWER. P(A and B) = P(A)*P(B) = 0.82*0.87. Use your calculator.
Solved, with full explanations.
|
|
|